OpenStudy (anonymous):

Solve 25x = 0.04. Round your answer to the nearest hundredth. Answer x = -2 x = -1 x = 1 x =

OpenStudy (anonymous):



OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

x=0.04/25=4/2500=1/625 none of your options are correct :( are you sure ?

OpenStudy (anonymous):

Yup only the options given @ghazi which one closect?

OpenStudy (ghazi):

the last one